OpenStudy (anonymous):

]Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

First change that equation into the standard y=mx +b form (where m is the slope, x and y are the coordinates (x,y) and b is the y intercept) You will have to substract x from both sides and get -4y = 4-x Then divide both sides by -4 get y by itself, getting you y= (-x+4)/-4 y=(1\4)x -1 The slope is (1/4, which is m) The y-intercept is (-1, which is b) Complete. Simple enough. :)

OpenStudy (anonymous):

put the equation into slope intercept form x - 4y = 4 x = 4y + 4 4y = x-4 y = 1/4*x - 1 The slope is then 1/4 and the intercept is -1.